OpenStudy (anonymous):

Express the triple integral y dv as integrals dzdydx and then dydzdx. E is the solid bounded by paraboloid z=x^2+y^2 and the planes z=4 and y=0 on the positive side of the y-axis. Sketch E.

OpenStudy (anonymous):

\[\int\limits_{-2}^{2}\int\limits_{-2}^{0}\int\limits_{0}^{x²+y²}y dzdydx\]

OpenStudy (anonymous):

Would it be easier to solve the other first?